# A novel efficient structure-preserving exponential integrator for Hamiltonian systems

### v1

@@ Abstract
We propose a linearly implicit structure-preserving numerical method for semilinear Hamiltonian systems with polynomial nonlinearities, combining Kahan’s method and exponential integrator. This approach efficiently balances computational cost, accuracy and the preservation of key geometric properties, including symmetry and near-preservation of energy. By requiring only the solution of a single linear system per time step, the proposed method offers significant computational advantages while comparing with the state-of-the-art symmetric energy-preserving exponential integrators. The stability, efficiency and long-term accuracy of the method are demonstrated through numerical experiments on systems such as the Henon-Heiles system, the Fermi-Pasta-Ulam system and the two-dimensional Zakharov-Kuznestov equation.
@@
structure-preserving, energy preservation, linearly implicit , exponential integrator, symmetric
1 Intorduction
@@
where $Q$ is a skew-symmetric matrix, $H$ is the Hamiltonian energy function containing a quadratic part and a higher-order part [6]. For such systems, the Hamiltonian remains invariant along the exact solution trajectory. Moreover, when $Q$ is a canonical skew-symmetric matrix, the system preserves the symplectic two-form. Maintaining these geometric structures is essential for the long-time numerical integration of such systems, as it ensures accuracy and stability [7]. To address these needs, various structure-preserving integrators have been proposed, including symmetric exponential integrators [4, 8], symplecticity-related exponential integrators [9, 10, 11, 12, 13], volume-preserving exponential integrator [14] and energy-preserving exponential integrators [15, 16]. Extensive studies have demonstrated that these specialized methods deliver significantly enhanced long-term performance compared to conventional high-order schemes.
@@
In recent years, energy-preserving exponential integrators have been widely studied due to their effectiveness in accurately capturing the energy dynamics of Hamiltonian systems [15, 17, 18, 19, 20, 21, 22, 23]. A prominent example is the exponential averaged vector field (EAVF) method [15] which combines the discrete gradient approach with exponential integrator. This hybrid strategy has inspired the development of a broad class of exponential energy-preserving schemes. The computation of such fully implicit methods requires solving nonlinear systems at each time step—a process that can be computationally demanding. As an alternative, linearly implicit structure-preserving methods have been explored. These methods offer similar stability as fully explicit schemes while requiring only the solution of a single linear system per time step, making them more computationally efficient. One class of such methods is based on the Scalar Auxiliary Variable (SAV) framework [24, 25, 18], which conserves a modified energy for semilinear Hamiltonian systems; however they are not symmetric. A number of further developments along these lines can be found in [12, 13, 21, 26, 27]. Another category involves multiple-point methods that portion out the nonlinearity [26, 8]. These methods have the advantage of preserving both a modified energy and symmetry, but they are inherently multistep-even for systems with quadratic vector fields-which may bring stability concerns when compared to single-step method [28].
In this work, we propose a linearly implicit structure-preserving method for semilinear Hamiltonian systems with polynomial nonlinearities, by combining Kahan’s method [29] with exponential integrators. Kahan’s method is employed due to several favorable properties: i) it is linearly implicit, requiring only the solution of a single linear system per time step, which improves computational efficiency compared to fully implicit methods; ii) it is symmetric, a key property for ensuring stable long-term integration; and iii) it preserves a modified energy and measure, geometric characteristics that are also satisfied by the exact flow of the original system. The proposed method inherits all advantageous properties of Kahan’s method except for the exact conservation of a modified energy and measure. However, we show that the energy error remains bounded and exhibits oscillatory behavior over time. Moreover, we derive an explicit analytical expression for the energy error, offering insights into the method’s long-term energy behavior. Through numerical experiments, we demonstrate that the proposed method achieves superior computational efficiency compared to both the fully implicit exponential AVF method and the linearly implicit method based on the polarized discrete gradient. Additionally, we provide a rigorous proof that the proposed method achieves second-order convergence.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the proposed linearly implicit structure-preserving exponential integrator and discuss its key properties, including symmetry, energy preservation and convergence. In Section 3, we present numerical experiments to demonstrate the stability, computational efficiency, and long-term behavior of the proposed method using different types of differential equations. Finally, conclusions and potential directions for future research are discussed in the last section.
@@ 2 The new linearly implicit structure-preserving symmetric method
Before proposing the new exponential integrator, we will introduce Kahan’s method. For a Hamiltonian system with a cubic Hamiltonian, Kahan’s method has the form
@@
For a semilinear system (1), we consider the following variation of constants formula on interval $[t_{0},t_{0}+h]$
$x(t_{0}+h)=e^{hA}x(t_{0})+h\int_{0}^{1}e^{(1-\tau)hA}f(x(t_{0}+\tau h))\mathrm{d}\tau.$ (3)
The key idea of constructing exponential integrator is to integrate the linear part of equation (3) exactly meanwhile finding an approximation to the integration on the right-hand side. For example, the exponential Euler method is obtained by approximating the nonlinear term by $f(x(t_{0}))$ [5], thereby having the form
@@
where $\phi(z):=\frac{e^{z}-1}{z}$; the exponential averaged vector field method is obtained by approximating the nonlinear term by the discrete gradient [15], possessing the form
[…]
In this work, we approximate the nonlinear term using Kahan’s method, leading to the new linearly implicit exponential integrator suitable for solving semilinear systems (1):
@@
We denote this newly derived scheme as the EKahan scheme, and its properties are described in detail below.
@@
The EKahan scheme (4) is symmetric.
@@
[…]
$x_{n}=e^{-Ah}x_{n+1}-h\phi(-Ah)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$ (5)
Multiplying both sides of (5) by $e^{hA}$, we get
$e^{hA}x_{n}=x_{n+1}-he^{hA}\phi(-Ah)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$ (6)
Using the identiy $\phi(-V)=e^{-V}\phi(V)$ for any matrix $V$, we can rewrite the above equation (6) as
$e^{hA}x_{n}=x_{n+1}-h\phi(hA)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$
This equation is equivalent to (4), confirming that the EKahan scheme (4) is symmetric. ∎
@@
The EKahan scheme (4) is second-order.
Proof.
By applying Taylor expansion to $f(\frac{x_{n}+x_{n+1}}{2})$ and $f(x_{n+1})$ at the point $x_{n}$ and using the fact that $f$ is quadratic, we can obtain
$-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})=f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n}).$
Therefore, EKahan scheme (4) can be rewritten as
$x_{n+1}$ $=e^{hA}x_{n}+h\phi(hA)[f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n})].$ (7)
Since the matrix exponential satisfies $e^{hA}=I+h\phi(hA)A$, we rewrite equation (7) as
$x_{n+1}$ $=(I+h\phi(hA)A)x_{n}+h\phi(hA)[f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n})]$
$=x_{n}+h\phi(hA)g(x_{n})+\frac{h}{2}\phi(hA)f^{\prime}(x_{n})(x_{n+1}-x_{n}),$ (8)
where the function $g(x)$ is the right-hand side of the semilinear equation (1), i.e., $g(x):=Ax+f(x)$. Supposing the matrix $I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n})$ is invertible, solving for $x_{n+1}-x_{n}$ from equation (2) gives
$x_{n+1}-x_{n}=h(I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n}))^{-1}\phi(hA)g(x_{n}).$
By $L^{\prime}H\hat{o}pital^{\prime}s\;rule$, we have $\lim\limits_{z\rightarrow 0}\frac{e^{z}-1}{z}=1$, so $\lim\limits_{h\rightarrow 0}\phi(hA)=I$ and
$\lim\limits_{h\rightarrow 0}(I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n}))^{-1}\phi(hA)g(x_{n})=g(x_{n}).$
Therefore, we have the following truncation for the EKahan scheme (7)
$x_{n+1}$ $=x_{n}+hg(x_{n})+O(h^{2}).$ (9)
By Taylor expansion, the exact solution of the semilinear Hamiltonian system (1) satisfies
$x(t_{n+1})=x(t_{n})+hg(x(t_{n}))+\frac{h^{2}}{2}(A+f^{\prime}(x(t_{n})))g(x(t_{n}))+O(h^{3}).$ (10)
Supposing $x_{n}$ is exactly solved and subtracting equation (10) from equation (9), we get
$x_{n+1}-x(t_{n+1})=O(h^{2}).$
Moreover, EKahan scheme (4) is symmetric from Lemma 1, confirming that this method is at least second-order. ∎
Theorem 1.
[…]
$\|f(x_{2})-f(x_{1})\|\leq L\|x_{2}-x_{1}\|.$
Then, the error bound
$\|x(t_{n})-x_{n}\|\leq e^{k_{2}T}(k_{1}\|x(t_{0})-x_{0}\|+k_{3}h^{2})$
holds in $0\leq nh\leq T$ for constant $k_{1},\;k_{2},\;k_{3}$.
Proof.
The assumptions about the operator $A$ guarantee the boundness of $\|\phi(hA)\|$ and $\|e^{jhA}\|$, $j=1,2,\dots$, and we denote the bound by $M$; see Lemma 2.4 from [5]. Since EKahan scheme (4) is second-order, we get the following truncation when inserting the exact solutions $x(t_{n})$ and $x(t_{n+1})$ to the numerical scheme:
$\begin{split}x(t_{n+1})=&e^{hA}x(t_{n})+h\phi(hA)[-\frac{1}{2}f(x(t_{n}))-\frac{1}{2}f(x(t_{n+1}))\\ &+2f(\frac{x(t_{n})+x(t_{n+1})}{2})]+O(h^{3}).\end{split}$ (11)
Denote the numerical error terms by $\varepsilon_{n}=x(t_{n})-x_{n}$ and
$\delta_{n+1,n}=$ $-\frac{1}{2}f(x(t_{n}))+2f(\frac{x(t_{n})+x(t_{n+1})}{2})-\frac{1}{2}f(x(t_{n+1}))$
$-\big{(}-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})\big{)}.$
Subtracting equation (4) from equation(11), we get the following error recursion equation
$\varepsilon_{n+1}=e^{hA}\varepsilon_{n}+h\phi(hA)\delta_{n+1,n}+O(h^{3}).$
This recursion equation gives
$\varepsilon_{n}=e^{nAh}\varepsilon_{0}+\sum\limits_{j=0}^{n-1}e^{jAh}h\phi(hA)\delta_{n-j,n-j-1}+\sum\limits_{j=0}^{n-1}e^{jAh}O(h^{3}).$ (12)
Since $f$ satisfies $Lipschitz\;condition$, we have
$\|\delta_{n-j,n-j-1}\|\leq\frac{3}{2}L\|\varepsilon_{n-j}\|+\frac{3}{2}L\|\varepsilon_{n-j-1}\|.$
Then we have
$\|\sum\limits_{j=0}^{n-1}e^{jAh}h\phi(hA)\delta_{n-j,n-j-1}\|\leq\frac{3}{2}hLM\|\varepsilon_{n}\|+3hLM^{2}\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|.$ (13)
Combing (12) and (13), we obtain
$\|\varepsilon_{n}\|\leq\|e^{TA}\|\|\varepsilon_{0}\|+\frac{3}{2}hLM\|\varepsilon_{n}\|+3hLM^{2}\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|+nMO(h^{3}).$ (14)
For small $h$ such that $1-\frac{3}{2}hLM$ is positive, suppose there exists a positive real number $k_{3}$ satisfying $\frac{nMO(h^{3})}{1-\frac{3}{2}hLM}\leq k_{3}h^{2}$. Denote by $k_{1}=\frac{\|e^{TA}\|\|\varepsilon_{0}\|}{1-\frac{3}{2}hLM}$ and $k_{2}=\frac{3LM^{2}}{1-\frac{3}{2}hLM}$. Then we can obtain the following estimation for the numerical error $\varepsilon_{n}$ from the above relation (14):
$\|\varepsilon_{n}\|$ $\leq k_{1}\|\varepsilon_{0}\|+k_{2}h\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|+k_{3}h^{2}$
$\leq e^{nk_{2}h}(k_{1}\|\varepsilon_{0}\|+k_{3}h^{2})$
$=e^{k_{2}T}(k_{1}\|\varepsilon_{0}\|+k_{3}h^{2}).$
The last inequality follows from the Gronwall’s inequality. ∎
Before presenting the second main theorem about the energy behaviour of EKahan method, we give the following lemma showing a property shared by a general exponential integrator based on the variational constant formula applied to the semilinear system (2). The proof of this lemma is inspired by Theorem 1 in [15].
Lemma 3.
For the semilinear system (2), any exponential integrator with the form

### v2

# arXiv:2506.07072v2[math.NA] 02 Mar 2026

@@ Abstract
We develop an efficient linearly implicit time-reversible exponential integrator for semilinear Hamiltonian systems with polynomial nonlinearities. The method combines the variation of constants formula with Kahan-type discretization of the nonlinear term. For quadratic vector fields, we prove second-order convergence in finite dimensions and derive an exact stepwise energy-increment identity implying a precise notion of near conservation of energy. For higher-degree polynomial Hamiltonians, we employ polarization to obtain a linearly implicit multistep exponential extension and discuss its well-posedness, stability and convergence; we further provide a stepwise energy characterization that quantifies its energy drift in terms of the numerical increments. Numerical experiments on the Hénon-Heiles system, a continuum Fermi-Pasta-Ulam model, and the two-dimensional Zakharov-Kuznetsov equation corroborate the theoretical accuracy and demonstrate competitive accuracy-cost tradeoffs compared with representative symmetric energy-preserving exponential integrators.
@@
Hamiltonian system, linearly implicit, exponential integrato, symmetric, near conservation of energy
1 Introduction
@@
i.e., $A=QM$ and $f(x)=Q\nabla U(x)$, where $Q$ is a skew-symmetric matrix, $H$ is the Hamiltonian energy function containing a quadratic part and a higher-degree part [5]. For such systems, the Hamiltonian remains invariant along the exact solution trajectory. Moreover, when $Q$ is a canonical skew-symmetric matrix, the system preserves the symplectic two-form. Maintaining these geometric structures is essential for the long-time numerical integration of such systems, as it ensures accuracy and stability [15]. To address these needs, various structure-preserving integrators have been proposed, including symmetric exponential integrators [4, 22], symplecticity-related exponential integrators [34, 33, 26, 17, 30], volume-preserving exponential integrators [32] and energy-preserving exponential integrators [23, 12]. Extensive studies have demonstrated that these specialized methods deliver significantly enhanced long-term performance compared to conventional high-order schemes.
@@
In recent years, energy-preserving exponential integrators have been widely studied due to their effectiveness in accurately capturing the energy dynamics of Hamiltonian systems [23, 29, 18, 27, 8, 19, 28, 21]. A prominent example is the exponential averaged vector field (EAVF) method [23] which combines the discrete gradient approach with exponential integrator. This hybrid strategy has inspired the development of a broad class of exponential energy-preserving schemes. Solving such fully implicit methods requires tackling nonlinear systems at each time step, which can be computationally demanding. As an alternative, linearly implicit structure-preserving methods have been explored. These methods offer similar stability to fully implicit schemes while requiring only the solution of a single linear system per time step, making them more computationally efficient. One class of such methods is based on the Scalar Auxiliary Variable (SAV) framework [31, 24, 18], which conserves a modified energy for semilinear Hamiltonian systems but lacks symmetry. Symmetry is also an important geometric property, as it enforces discrete time-reversibility, yields an even-power error expansion, and often improves long-time behavior [15]. A number of further developments along these lines can be found in [17, 30, 19, 35, 13]. Another category involves multiple-point methods that portion out the nonlinearity [25, 22]. These methods have the advantage of preserving both a modified energy and symmetry, but they may suffer from stability concerns when compared to Kahan methods [9], and a systematic convergence theory for linearly implicit time-reversible exponential integrators, especially in the higher-degree polynomial setting, is still limited [22]. These observations motivate the design of stable time-reversible and linearly implicit exponential integrators, together with a rigorous convergence analysis. We therefore incorporate the Kahan discretization for the nonlinear term, since it is linearly implicit, time-reversible and known to preserve a modified energy and a modified measure [6, 7].
To treat the stiff linear dynamics in (1), we embed Kahan update into the variation of constants exponential framework, leading to symmetric schemes that require only a linear solve per step and admit a quantitative near conservation of energy. This coupling is, however, nontrivial: the exponential update changes the algebraic structure of the original Kahan discretization, so its classical modified-energy preservation mechanism does not carry over in general, and the energy conservation analysis must be rebuilt. Our main contributions in this work are as follows. First, for quadratic vector fields (cubic Hamiltonians), we obtain a one-step linearly implicit symmetric exponential scheme and prove second-order convergence on finite time intervals. Second, for higher-degree polynomial Hamiltonians, we derive a linearly implicit time-reversible multistep exponential extension via polarization and establish solvability and stability analysis. Third, we derive explicit stepwise energy increment identities, which yield computable energy-drift estimates in terms of numerical increments and explain the favorable long-time behavior observed in our experiments.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the proposed linearly implicit exponential integrator for cubic Hamiltonians, discuss its key properties, including symmetry, convergence, near preservation of energy, and its generalization to systems with higher-degree polynomial Hamiltonians. In Section 3, we present numerical experiments to demonstrate the stability, computational efficiency, and long-term behavior of the proposed method using different types of differential equations. Finally, conclusions and potential directions for future research are discussed in the last section.
@@ 2 The new linearly implicit structure-preserving symmetric method
Before proposing the new exponential integrator, we will introduce Kahan’s method and exponential integrators since our new scheme is created based on them. For system (2) with a cubic Hamiltonian, Kahan’s method has the form
@@
For a semilinear system (1), we consider the following variation of constants formula on interval $t_{0},t_{0}+h]$
$x(t_{0}+h)=e^{hA}x(t_{0})+h\int_{0}^{1}e^{(1-\tau)hA}f(x(t_{0}+\tau h))\mathrm{d}\tau.$
The key idea of constructing exponential integrator is to integrate the linear part exactly meanwhile finding an approximation to the integration of the nonlinear part. For example, the exponential Euler method is obtained by approximating the nonlinear term by $f(x(t_{0}))$ [16], thereby having the form
@@
where $\phi(z):=\frac{e^{z}-1}{z}$; the exponential averaged vector field (EAVF) method is obtained by approximating the nonlinear term by the discrete gradient [23], possessing the form
$x_{n+1}=e^{hA}x_{n}+h\phi(hA)\int_{0}^{1}f(\xi x_{n}+(1-\xi)x_{n+1})d\xi.$
2.1 The one-step EKahan scheme for cubic Hamiltonians
For system (2) with a cubic Hamiltonian, we propose a linearly implicit exponential integrator by approximating the nonlinear term using the above mentioned Kahan discretization, denoted by EKahan
@@ 2.1 The one-step EKahan scheme for cubic Hamiltonians
Before presenting the properties of scheme (3), we give the following notations and assumptions:
Norms and growth constants
Let $\|\cdot\|$ be an induced matrix norm and the associated vector norm. For a fixed final time $T>0$, define
$M_{E}=\max_{0\leq t\leq T}\|e^{tA}\|,\qquad M_{\phi}=\max_{0\leq t\leq T}\|\phi(tA)\|.$
Lipschitz assumptions
For system (2), we assume that there exists a convex bounded set $\mathcal{D}\subset\mathbb{R}^{d}$ such that both the exact solution $\{x(t):t\in[0,T]\}$ and the numerical solution $\{x_{n}:0\leq nh\leq T\}$ remain in $\mathcal{D}$, and that $f$ is Lipschitz on $\mathcal{D}$ with constant $L$.
@@
Assuming that $f$ is a quadratic vector field and that the exact solution of (1) is sufficiently smooth on $t_{n},t_{n+1}]$, then the EKahan scheme (3) has local truncation error $O(h^{3})$.
@@
For quadratic $f$, we have the exact Taylor formulas
$f(x_{n+1})$ $=f(x_{n})+f^{\prime}(x_{n})\Delta_{x}+\tfrac{1}{2}f^{\prime\prime}(x_{n})(\Delta_{x},\Delta_{x}),$
$f\left(\tfrac{x_{n}+x_{n+1}}{2}\right)$ $=f(x_{n})+\tfrac{1}{2}f^{\prime}(x_{n})\Delta_{x}+\tfrac{1}{8}f^{\prime\prime}(x_{n})(\Delta_{x},\Delta_{x}),$
where $\Delta_{x}=x_{n+1}-x_{n}$ and $f^{\prime\prime}$ is constant. Hence, we have
$\tfrac{1}{2}f(x_{n})+2f\!\left(\tfrac{x_{n}+x_{n+1}}{2}\right)-\tfrac{1}{2}f(x_{n+1})=f(x_{n})+\tfrac{1}{2}f^{\prime}(x_{n})\Delta_{x}.$ (4)
Using (4), the EKahan scheme (3) can be rewritten as
$x_{n+1}=e^{hA}x_{n}+h\phi(hA)\Bigl(f(x_{n})+\tfrac{1}{2}f^{\prime}(x_{n})\Delta_{x}\Bigr).$ (5)
Since $e^{hA}=I+h\phi(hA)A$ and $A$ commutes with $\phi(hA)$, (5) becomes
$x_{n+1}=x_{n}+h\phi(hA)g(x_{n})+\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})\Delta_{x},\quad\text{with}\quad g(x)=Ax+f(x),$
and therefore we have
$\Bigl(I-\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})\Bigr)\Delta_{x}=h\phi(hA)g(x_{n}).$ (6)
For sufficiently small $h$, the matrix in (6) is invertible, and we may write
$\Delta_{x}=h\Bigl(I-\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})\Bigr)^{-1}\phi(hA)g(x_{n}).$ (7)
Using the series expansion $\phi(hA)=I+\tfrac{h}{2}A+O(h^{2})$ and the Neumann expansion
$\Bigl(I-\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})\Bigr)^{-1}=I+\tfrac{h}{2}\phi(hA)f^{\prime}(x_{n})+O(h^{2}),$
we obtain from (7) that
$\Delta_{x}=h\,g(x_{n})+\tfrac{h^{2}}{2}\bigl(A+f^{\prime}(x_{n})\bigr)g(x_{n})+O(h^{3}).$
Consequently, we have
$x_{n+1}=x_{n}+h\,g(x_{n})+\tfrac{h^{2}}{2}\bigl(A+f^{\prime}(x_{n})\bigr)g(x_{n})+O(h^{3}).$ (8)
On the other hand, the exact solution satisfies
$x(t_{n+1})=x(t_{n})+h\,g(x(t_{n}))+\tfrac{h^{2}}{2}g^{\prime}(x(t_{n}))g(x(t_{n}))+O(h^{3}),$ (9)
where $g^{\prime}(x)=A+f^{\prime}(x)$. Taking $x_{n}=x(t_{n})$ and comparing (8) with (9) yields $x_{n+1}-x(t_{n+1})=O(h^{3})$.
Theorem 1.
Assume $f$ satisfies the standard Lipschitz condition and the local truncation error satisfies $\|\tau_{n+1}\|\leq C_{\rm loc}h^{3}$ for all $0\leq nh\leq T$, which is true by Lemma 1. Then there exist constants $C_{1},C_{2}>0$, depending only on $\|A\|$, $L$, $T$, $M_{E}$, $M_{\phi}$ and $C_{\rm loc}$, such that
$\|x(t_{n})-x_{n}\|\leq e^{C_{1}T}\Bigl(\|x(t_{0})-x_{0}\|+C_{2}h^{2}\Bigr),\quad 0\leq nh\leq T,$
if the step size satisfies
$h\leq h_{0}:=\frac{1}{3LM_{\phi}}.$ (10)
In particular, if $x_{0}=x(t_{0})$, then $\max\limits_{0\leq nh\leq T}\|x(t_{n})-x_{n}\|=O(h^{2})$.
Proof.
Set $\varepsilon_{n}=x(t_{n})-x_{n}$ and $\Psi(u,v)=-\tfrac{1}{2}f(u)+2f\!\Bigl(\tfrac{u+v}{2}\Bigr)-\tfrac{1}{2}f(v)$. Inserting the exact solution into (3) yields
$x(t_{n+1})=e^{hA}x(t_{n})+h\phi(hA)\Psi(x(t_{n}),x(t_{n+1}))+\tau_{n+1},\quad\|\tau_{n+1}\|\leq C_{\rm loc}h^{3}.$ (11)
Subtracting (3) from (11) gives the error recursion
$\varepsilon_{n+1}=e^{hA}\varepsilon_{n}+h\phi(hA)\delta_{n+1,n}+\tau_{n+1},$ (12)
where $\delta_{n+1,n}=\Psi(x(t_{n}),x(t_{n+1}))-\Psi(x_{n},x_{n+1})$. Since $\mathcal{D}$ is convex and $f$ is Lipschitz on $\mathcal{D}$, we have
$\|\delta_{n+1,n}\|$ $\leq\tfrac{1}{2}\|f(x(t_{n}))-f(x_{n})\|+2\Bigl\|f\!\Bigl(\tfrac{x(t_{n})+x(t_{n+1})}{2}\Bigr)-f\!\Bigl(\tfrac{x_{n}+x_{n+1}}{2}\Bigr)\Bigr\|$
$+\tfrac{1}{2}\|f(x(t_{n+1}))-f(x_{n+1})\|$
$\leq\tfrac{3}{2}L(\|\varepsilon_{n}\|+\|\varepsilon_{n+1}\|).$
Taking norms in (12) and using $\|\phi(hA)\|\leq M_{\phi}$ yields
$\|\varepsilon_{n+1}\|\leq\|e^{hA}\|\,\|\varepsilon_{n}\|+\tfrac{3}{2}LhM_{\phi}\bigl(\|\varepsilon_{n}\|+\|\varepsilon_{n+1}\|\bigr)+C_{\rm loc}h^{3}.$ (13)
Denoting by $\alpha=\tfrac{3}{2}LhM_{\phi}$, moving the term $\alpha\|\varepsilon_{n+1}\|$ to the left in (13) gives
$(1-\alpha)\|\varepsilon_{n+1}\|\leq(\|e^{hA}\|+\alpha)\|\varepsilon_{n}\|+C_{\rm loc}h^{3},$
and hence we have
$\|\varepsilon_{n+1}\|\leq q(h)\,\|\varepsilon_{n}\|+\frac{C_{\rm loc}}{1-\alpha}h^{3},\quad\text{with}\quad q(h)=\frac{\|e^{hA}\|+\alpha}{1-\alpha}.$ (14)
Using $e^{hA}=I+\int_{0}^{h}Ae^{sA}\mathrm{d}s$ and $\|e^{sA}\|\leq M_{E}$ for $0\leq s\leq h$, we obtain
$\|e^{hA}\|\leq 1+h\|A\|M_{E}.$
Moreover, under (10) we have $\alpha\leq\tfrac{1}{2}$, so that $(1-\alpha)^{-1}\leq 1+2\alpha$ and $2\alpha^{2}\leq\alpha$. Therefore, we get
$q(h)=\frac{\|e^{hA}\|+\alpha}{1-\alpha}\leq(\|e^{hA}\|+\alpha)(1+2\alpha)\leq 1+Ch$
with $C=\|A\|M_{E}+3(1+M_{E})LM_{\phi}$. In addition, since $\alpha\leq\tfrac{1}{2}$ we have $(1-\alpha)^{-1}\leq 2$, and thus
$\frac{C_{\rm loc}}{1-\alpha}h^{3}\leq 2C_{\rm loc}h^{3}.$
Substituting these bounds into (14) gives
$\|\varepsilon_{n+1}\|\leq(1+Ch)\|\varepsilon_{n}\|+2C_{\rm loc}h^{3},\qquad 0\leq nh\leq T,\ \ h\leq h_{0}.$ (15)
Applying the discrete Gronwall inequality to (15) yields, for $t_{n}=nh\leq T$,
$\|\varepsilon_{n}\|\leq(1+Ch)^{n}\|\varepsilon_{0}\|+2C_{\rm loc}h^{3}\sum_{j=0}^{n-1}(1+Ch)^{j}.$
Using $(1+Ch)^{n}\leq e^{CT}$ for $nh\leq T$ and
$\sum_{j=0}^{n-1}(1+Ch)^{j}=\frac{(1+Ch)^{n}-1}{Ch}\leq\frac{(1+Ch)^{n}}{Ch}\leq\frac{e^{CT}}{Ch},$
we obtain
$\|\varepsilon_{n}\|\leq e^{CT}\|\varepsilon_{0}\|+\frac{2C_{\rm loc}}{C}e^{CT}h^{2},$
which proves the stated estimate with $C_{1}=C$ and $C_{2}=2C_{\rm loc}/C$. ∎
Before presenting the second main theorem about the energy behaviour of EKahan method, we give the following lemma showing a property shared by a general exponential integrator based on the variation of constant formula applied to the semilinear system (2). The proof of this lemma is inspired by Theorem 1 in [23].
@@
For system (2), any exponential integrator of the form